Sigma n=1 to infinity (-1)^n sin(1/n)
1
Expert's answer
2021-09-27T16:22:48-0400

This is an alternating series. We shall use alternating series test to check if it converges or not.

"\lim_{n\to \infty}\sin(\frac{1}{n})=\sin(0)=0"

Hence, the series "\sum_{n=1}^{\infty} (-1)^n \sin(\frac{1}{n})" is convergent by the alternating series test.